OpenStudy (anonymous):

5x+4=3(2x-1)-x+7 Solve for X

OpenStudy (anonymous):

Distribute the 3.

OpenStudy (anonymous):

5x+4=6x-3-x+7 5x+4=5x+4 5x=5x x=1

OpenStudy (anonymous):

@Edutopia so then it would be 5x+4=6x-1-x+7

jimthompson5910 (jim_thompson5910):

3(2x-1) = 3*2x - 3*1 = 6x - 3 So 3(2x-1) = 6x-3

jimthompson5910 (jim_thompson5910):

So after distributing, it goes from 5x+4=3(2x-1)-x+7 to 5x+4=6x-3-x+7

OpenStudy (anonymous):

@jim_thompson5910 ok, so what do i do next ?

jimthompson5910 (jim_thompson5910):

Combine like terms on the right side